OpenStudy (hawkfalcon):

Find cos(theta) if tan(theta) = (-2/3) and sin(theta) > 0

OpenStudy (hawkfalcon):

Find \[\cos \theta if \tan \theta = -2/3 and \sin \theta >0\]

OpenStudy (anonymous):

Here, you know that the tangent is opp/adjacent. Since the sine is positive, it must be in either the first or second quadrant right?

OpenStudy (hawkfalcon):

Yes.

OpenStudy (anonymous):

So the opposite leg must be positive, therefore since the tangent is negative the adjacent leg must be the negative one. So what quadrant is that?

OpenStudy (hawkfalcon):

Oh so you know o and a, you just need to find the h, and you can solve it :P

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Lets see what ya got when ya get it.

OpenStudy (hawkfalcon):

Thats 2nd quad

OpenStudy (hawkfalcon):

Okay:)

OpenStudy (anonymous):

Yep.

OpenStudy (hawkfalcon):

\[\frac{ -3 }{ \sqrt{13}}\]

OpenStudy (anonymous):

Thats it.

OpenStudy (hawkfalcon):

:D thank you!

OpenStudy (anonymous):

No prob.

OpenStudy (anonymous):

\[\frac{ x \sqrt{\left| x ^{2}-4 \right|} }{ x ^{2}-4 }-1>0\]

OpenStudy (hawkfalcon):

@Yasir_se wtf?